OpenStudy (anonymous):

The midpoint of an has coordinates (2,3) and A is at the orgin. How can you find the coordinates of B? Give the coordinates of b.

OpenStudy (anonymous):

2 = (0 +x ) / 2 x = 4 3 = (0+y) / 2 y = 6 Co-ordinates of B : (4,6) Hope this helps :)